OpenStudy (anonymous):

Solve and Check. 5x^2-13x=0

OpenStudy (anonymous):

x(5x-13) = 0

OpenStudy (anonymous):

x(5x - 13) = 0 x = 0 or x = 13/5

OpenStudy (anonymous):

@PECKCE

OpenStudy (anonymous):

Thank you.